OpenStudy (anonymous):

4w^2-(3z+1)^2 Use difference of squares pattern to factor. Any suggestions? I have been beaten down with a few of my problems this week.

OpenStudy (anonymous):

(2w+3z+1)(2w-3z-1) this what your asking for?

OpenStudy (anonymous):

Correct, but how in the world do I get there? I have missed some step.

OpenStudy (anonymous):

just reverse foil.

OpenStudy (kainui):

Since you can see that this very general, kind of abstract thing is true: (a+b)(a-b)=a^2+ab-ab-b^2=a^2-b^2 then if you can just find something squared minus something else squared, you are really just following the pattern. So in this case, a=(2w) and b=(3z+1). Does this make sense more or less? Think about it a little and see if you can sort of use the relation I showed above to get you from one side to the other.

OpenStudy (anonymous):

Thank you both! I appreciate the help!

OpenStudy (anonymous):

You both had best response in my opinion. Is there a math problem to figure out how to give you both the thumbs up?

OpenStudy (kainui):

Raise both your thumbs, take away the medal from him, then give it to me. Problem solved!